OpenStudy (anonymous):

find the product -9y^2(y^4-7y^3-y^2+2y+5) can someone please help me to solve this please

OpenStudy (callisto):

\[-9y^2(y^4-7y^3-y^2+2y+5)\]\[= y^4 (-9y^2) -7y^3(-9y^2) -y^2 (-9y^2) +2y(-9y^2)+5 (-9y^2)\]\[= -9^{4+2} -7(-9) y^{3+2} - (-9)y^{2+2} + 2(-9)y^{1+2} -9y^2\]Can you simplify it?

OpenStudy (anonymous):

-9y+63y^5+9y^4-18y^3-45y^2 ?

OpenStudy (callisto):

Oh... I made a mistake in my previous post :| The last term there should be -45y^2 , which you've got it correct:) The first term -> not correct The rest are correct :)

OpenStudy (anonymous):

so this is the correct answer

OpenStudy (callisto):

Nope ._. The first term is not -9y it's -9y^{4+2} =?

OpenStudy (anonymous):

-9y^6

OpenStudy (callisto):

Yes :) So the final answer is?

OpenStudy (anonymous):

-9y^6+63y^5(y^4-18y^3-45y^2 =

OpenStudy (callisto):

Hmm... It is: -9y^6+63y^5+9y^4-18y^3-45y^2

OpenStudy (anonymous):

yes that is what i meant to put, sorry typo

OpenStudy (anonymous):

can you help me with another one please

OpenStudy (callisto):

No problem :)

OpenStudy (anonymous):

can you

OpenStudy (callisto):

Of course :)